A voltage source š‘£(š‘”) = 240 sin(314.16š‘” āˆ’ 20Ā°)š‘‰ is connected to a load having an

impedance of Z. The resulting current through the load is š‘–(š‘”) = 15 sin(314.16š‘” + 22.5Ā°)A.

Determine the circuit power factor and identify the elements that constitute the load, given

that Z comprises of only two elements connected in series.


1
Expert's answer
2021-05-31T06:19:56-0400

i leads the v (voltage). This is possible for the capacitive loads.

Therefore given load is resistor and capacitor load.

"R=|Z|cos \phi = 16 cos -42.5=-10.81 \Omega"

"R=|Z|sin \phi = 16 sin 42.5=11.81 \Omega"

"Z=R+jX=(11.81-j10.81) \Omega"

The power factor is "cos \phi =cos 42.5 =0.74" leading

"X_c= \frac{1}{wc} \implies c= \frac{1}{X_cw}=\frac{1}{10.81*314.6}=0.00029404"

"C= 294.04*10^{-6} F"